OpenStudy (anonymous):

A man invests Rs.5000 for 3 years at 5% compound interest reckoned yearly. Income tax @ 20% on the interest earned is deducted at the end of the each year. Find the amount at the end of the third year

OpenStudy (anonymous):

After one year, Total amount=5000+5000*1*5/100 - 20/100 *5000*1*5/100 = 5200 After 2 year, Total amount = 5200 + 5200 * 1*5/100 - 20/100 *5000*1*5/100 =5408 After 3 year, total amount =5408+5408*1*5/100-20/100 *5408*1*5/100=5624.32

OpenStudy (anonymous):

tthnxxx buddy!!